OpenStudy (anonymous):

A rocket sled is tested at "5 g" (5 times the acceleration due to gravity). If the sled starts from rest at position do= 0.00, how long does it take to travel 441 meters?

OpenStudy (anonymous):

Since we know the acceleration and distance, and are looking for the time, we can use a useful kinematics formula: \[d=v_o t + \frac{ 1 }{ 2 }at^2\] Substituting in our values for the distance and acceleration, and since it starts from rest we can plug in 0 for the initial velocity: \[441m=\frac{ 1 }{ 2 }(49 m/s^2)t^2\] Solving for the time, we find: \[882m=(49m/s^2)t^2\] \[18 s^2 = t^2\] \[\sqrt{18}s =t\]